# DS communications: Trends, Challenges, and Solutions

### v3

@@ Abstract
With a large number of deep space (DS) missions anticipated by the end of this decade, reliable and high capacity DS communications are needed more than ever. Nevertheless, existing technologies are far from meeting such a goal. Improving current systems does not only require engineering leadership, but also, very crucially, investigating potential technologies that overcome the unique challenges of ultra-long DS links. To the best of our knowledge, there has not been any comprehensive surveys of DS communications technologies over the last decade. Free space optical (FSO) is an emerging DS technology, proven to acquire lower communications systems size, weight and power (SWaP) and achieve a very high capacity compared to its counterpart radio frequency (RF), the currently used DS technology. In this survey, we discuss the pros and cons of deep space optical communications (DSOC) and review physical and networking characteristics. Furthermore, we provide, for the very first time, thoughtful discussions about implementing orbital angular momentum (OAM) and quantum communications (QC) for DS. We elaborate on how these technologies among other field advances, including interplanetary network, and RF/FSO systems improve reliability, capacity, and security. This paper provides a holistic survey in DSOC technologies gathering 200+ fragmented literature and including novel perspectives aiming to setting the stage for more developments in the field.
@@
Deep space optical communications (DSOC), delay tolerant networking (DTN), detection technologies, interplanetary network (IPN), link capacity, modulation and coding schemes, orbital angular momentum (OAM), quantum communications (QC), radio frequency (RF)/free space optical (FSO) systems, networking protocols.
@@ I Introduction
Space exploration offers an opportunity to understand the universe, resulting in endless benefits for human life diverse aspects such as technologies, transportation, environment, etc [1]. There is an increased investment from several worldwide governmental space agencies [2] as well as private companies [3] aiming to expand human presence in the Solar system [4, 5]. A number of missions have been recently conducted, e.g. [6, 7, 8] to understand Mars atmosphere and search for life there. Many other missions are scheduled for the next several years, such as Rosalind Franklin [9]. There are also plans to send crews to Mars as early as 2026 [10].
@@ I-A Motivation
Space exploration is built upon reliable broadband interplanetary communications [11], which are only possible through developing dedicated technologies suitable for facing the unique deep space (DS) communications challenges. Unlike Earth and near-Earth, DS communications suffer from non-guaranteed line-of-sight, ultra-long time-changing distances, variability of Sun-Earth-probe angles, and the coronal or Solar wind plasma, referred to as Solar scintillation [12]. These challenges tremendously deteriorate the quality of the links and hence, limit capacity [13]. Current missions to Mars, for example, have only a few megabits per second communications rates at the minimum distance [14, 15, 16]. Fig. 1 shows a timeline for a number of Mars orbiters, landers, and rovers along with their data rates [7, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 8]. Capacity drops by orders of magnitudes as the distance increases [7]. Moreover, higher attenuation is seen beyond Mars missions. For example, the Voyager 1 spacecraft, which is located approximately at 24.1 billion km distance as of October 2023, operates at a maximum data rate of few kilo bits per second for less than a minute, only a few times a week [27].
Figure 1: Mars missions timeline with their communications links to Earth data rates.
To build high capacity DS communications, research efforts have been recently devoted to exploring the potentials of free space optical (FSO) communications [28, 13, 29, 30, 31, 32]. Compared to radio frequency (RF), FSO links enjoy larger bandwidth, smaller apertures at final stations, much-improved security, less transmit power and higher immunity against interference. Although most recent DS Mars missions are RF-based [16], many projects have already adopted FSO for Moon, and Mars communications with plans to expand to other planets such as Jupiter [33, 34, 35, 36, 37]. In addition, a number of implementation-based studies have been done to evaluate the technology transition, e.g. [30, 38]. A tremendous work is yet to be done to make FSO technology ready for future missions [39].
Adopting FSO in DS will play an important role in implementing interplanetary broadband networks. Achieving such a goal, though, will depend on the reliability, effectiveness, and efficiency of so many other technologies, including modulation, coding, detection, and networking protocols. The criteria for selecting DS communications schemes differ tremendously from those for terrestrial, airborne, and even near-Earth communications. In this survey, we highlight the various DS optical communications (DSOC) design criteria and present the latest research and innovations in the area.
@@ I-B Related Surveys
There are few surveys available on DS communications [29, 31, 36, 40, 30, 41, 42, 43, 44], most of which are outdated and/or do not address the field comprehensively. Although [29] covers various aspects of DSOC, an updated survey that reflects the last decade field demands, challenges, and advances is needed. [31] is another survey that covers optical communications in DS, but only partially. Biswas et al. in [36], on the other hand, focus merely on the progress made and status of DSOC national aeronautics and space administration (NASA) projects. In the same direction, [40] overviews mainly laser-based DS demonstration projects and future missions. [30], however, provides only studies about DS systems architects and discussions about the transition from RF to optical technology. In addition, while [41] focuses only on DS networking aspects, [42] covers mainly some aspects of RF-based DS communications systems. Although [43] addresses FSO and RF challenges, and discusses DS systems implementations and demonstrations, it misses presenting research discussions and future perspectives in the field. Finally, [44] gives an inspiring relatively-recent overview on DS networking and implementations, however, it lacks any further details in the other aspects, including physical characteristics and field advances.
Despite of their valuable contributions, none of the work above covers the DS communications holistically. In addition, most of these surveys are not recent. Our work, however, aims to cover the key aspects of DSOC and add the missing areas from the aforementioned related work while discussing the latest advances in the field, as we demonstrate in Table I.
@@ I-C Contributions
In this survey, we cover existing DSOC systems' limitations, recommended technologies to cope with them, and challenges facing their adoption. The contributions of this paper relative to the literature are summarized as follows.
@@
    We cover major challenges of DS communications. We also discuss DSOC promises, challenges, and potential solutions and compare it with existing RF technology.
@@
    We survey the widely used modulation and coding schemes, and the opportunities for enhancing performance by fusing them with emerging detection techniques.
@@
    We review recommended networking architecture models and present the key characteristics of promising protocols and identify related open research directions.
@@
    We present an overview of the interplanetary network (IPN) and study the visions for its future.
@@
    To the best of our knowledge, we provide the first thoughtful discussion about implementing orbital angular momentum (OAM) for DS and identify challenges and potential solutions.
@@
    We discuss the advantages, efforts, and challenges for implementing QC and partnering RF and FSO technologies in DS.
@@
TABLE I: DS COMMUNICATIONS SURVEYS COMPARISON
@@
Overview Missions timeline $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$ $\times$ $\surd$ $\surd$
Demonstrations $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Existing systems $\times$ $\surd$ $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Limitations $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Implementations $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\surd$
Future prospective $\surd$ $\surd$ $\surd$ $\surd$ $\times$ $\surd$ $\surd$ $\times$ $\surd$ $\surd$
Basics Link engineering $\surd$ $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$
Overall challenges $\surd$ $\surd$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\surd$ $\times$ $\surd$
Enabling technologies Protocols $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\surd$ $\surd$
Modulation $\surd$ $\surd$ $\times$ $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\times$ $\surd$
Coding $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\surd$
Receiver technologies $\surd$ $\surd$ $\times$ $\surd$ $\surd$ $\times$ $\surd$ $\surd$ $\surd$ $\surd$
Performance analysis Capacity $\surd$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$
Advances OAM $\times$ $\times$ $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\surd$
IPN $\times$ $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$
Quantum $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$
RF/FSO $\times$ $\surd$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\surd$
@@ I-D Organization
The remainder of the paper is structured as follows. Section II outlines DSOC advantages over RF, and implementation efforts and challenges. Section III analyzes DS links physical characteristics and discusses related research directions. While Section IV characterizes DSOC capacity limits, Section V surveys DS networking protocols and related research efforts. Section VI investigates on state-of-the-art and challenges of DSOC advances including, IPN, OAM, QC, and RF and FSO integration. Finally, Section VII concludes the survey.
@@ II DSOC Advantages and Challenges
Optical communications promise tens to a hundredfold increase in data rates, thanks to the high available bandwidth in the license-free spectrum in which the technology operates [31, 45]. Such a capacity improvement has already been proven in demonstrations [46]. To illustrate further the gain of using DSOC compared to RF, we, based on [39], define the figure of merit (FoM) metric as
@@
where $Dis$ is the end-to-end DS link distance in the astronomical unit (AU). $R$, $P$, and $Dia$ are the data rate in bits per second, transmit power in Watts, and receiver diameter size in meters, respectively. As it accounts for space terminal performance, measured through data rate, and cost, reflected through receiver aperture size and communications power, at any distance, the FoM metric provides a direct comparison between the RF and FSO technologies at the system level.
Based on projections available in [14, 40, 46], Fig. 2 provides a comparison of RF and FSO Lunar and Mars systems. Unsurprisingly, a Lunar FSO system represents more than 15 dB gain over the RF counterpart. The improvement seems to be even way more substantial over Mars distances. However, implementing FSO in DS communications faces a number of unique challenges.
Below we discuss different challenges facing DS in general at the link level, the DSOC link budget, and the optical communications challenges.
@@ II-A DS Communications Challenges
Ensuring high capacity and reliable communications in DS faces the challenges listed below.
  • Ultra-long, which based on the International Telecommunications Union exceeds 2 million kilometers, and time-changing distance between Earth and probes (satellites in other planets' orbiters, rovers, etc.).
@@
The communications with targets in DS require sending signals over very long distances, leading to strong power dissipation and hence a huge signal attenuation. Even with relatively large gain transmit and receive antennas, the received power can still be very weak. Also, the long-distance links result in an inevitable significant latency that prevents real-time communications and control. For example, a two-way communications with Mars may undergo 22 minutes transmission delay.
Furthermore, the Solar flares are shown to make communications signals weaker and noisier [29]. Besides, Solar scintillation and turbulence lead to performance degradation. It is important to carefully account for the near-Sun effect on the signal strength to compensate for any losses while preventing any damage to receiver circuits. Also, due to the movement of Earth and probes, links capacity get hugely affected. The Sun may completely block communications path.
In addition, SWaP constraints are of a major concern not only for efficient spacecraft design but also for communications performance improvement. Designing a space mission with ultra-lightweight resources is thought of as a way to utilize a better communications performance [29]. There are technologies such as Gossamer that has the premise to design spacecraft structures with thin, low-modulus, and low-mass materials. An example of Gossamer technology application is the inflatable antennas that could have an areal density of a fraction of a kilogram per square meter [47]. Also, in the direction of optical communications, the development of ultra-lightweight telescopes is realized, such as Cassegrain telescope that has composite mirrors, with areal density in the range of 1–10 kg/m2 [48].
Lastly, existing terrestrial technologies, especially at the physical and protocol levels, can not be adopted directly in DS. Digital modulations schemes are preferred, in DS RF-based communications for example, as they are simpler in implementation, less lossy, and more power and bandwidth efficient. Nevertheless, amplitude modulations that can not be generated with nonlinear amplifiers are not suitable for DS as working in saturation mode is needed to maximize power efficiency [49]. Also, in the networking, commercial terrestrial network architecture, for example transmission control protocol (TCP)/internet protocol (IP), are not efficient to ensure data delivery in the DS time-changing environments [50].

### v4

@@ Abstract
With a large number of deep space (DS) missions anticipated by the end of this decade, reliable and high-capacity DS communications are needed more than ever. Nevertheless, existing technologies are far from meeting such a goal. Improving current systems does not only require engineering leadership, but also, very crucially, investigating potential technologies that overcome the unique challenges of ultra-long DS links. To the best of our knowledge, there has not been any comprehensive surveys of DS communications technologies over the last decade. Free space optical (FSO) is an emerging DS technology, proven to acquire lower communications systems size weight and power (SWaP) and achieve a very high capacity compared to its counterpart radio frequency (RF), the currently used DS technology. In this survey, we discuss the pros and cons of deep space optical communications (DSOC) and review their physical and networking characteristics. Furthermore, we provide, for the first time, thoughtful discussions about implementing orbital angular momentum (OAM) and quantum communications (QC) for DS. We elaborate on how these technologies among other field advances including interplanetary network (IPN) and RF/FSO systems improve reliability, capacity, and security. This paper provides a holistic survey of DSOC technologies gathering 247 fragmented pieces of literature and including novel perspectives aiming to set the stage for more developments in the field.
@@
Deep space optical communications (DSOC), delay tolerant networking (DTN), detection technologies, interplanetary network (IPN), link capacity, modulation and coding schemes, networking protocols, orbital angular momentum (OAM), quantum communications (QC), radio frequency (RF)/free space optical (FSO) systems.
@@ I Introduction
Space exploration offers an opportunity to understand the universe, resulting in endless benefits for human life in diverse aspects such as technologies, transportation, environment, etc. [1]. There is an increased investment from several worldwide governmental space agencies [2] as well as private companies [3] aiming to expand human presence in the Solar system [4, 5]. Several missions have been recently conducted, e.g. [6, 7, 8], to understand Mars’s atmosphere and search for life there. Many other missions are scheduled for the next several years, such as the Rosalind Franklin explorer program [9]. There are also plans to send crews to Mars as early as 2029 [10].
@@ I-A Motivation
Space exploration is built upon reliable broadband interplanetary communications [11], which are only possible through developing dedicated technologies suitable for facing the unique challenges of deep space (DS) communications. Unlike Earth and near-Earth, DS communications suffer from non-guaranteed line-of-sight, ultra-long time-changing distances, variability of Sun-Earth-probe angles, and the coronal or Solar wind plasma, referred to as Solar scintillation [12]. These challenges tremendously deteriorate the quality of the links and hence, limit communications capacity [13]. Current missions to Mars, for example, have only a few megabits per second communications rates at the minimum distance [14, 15, 16]. Fig. 1 shows a timeline for several Mars orbiters, landers, and rovers along with their data rates [7, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 8]. It should be noted that higher attenuation is seen beyond Mars missions and the capacity drops by orders of magnitudes as the distance increases [7]. For example, the Voyager 1 spacecraft, which is located approximately 24.34 billion km from Earth as of May 2024, operates at a maximum data rate of a few kilo bits per second for less than a minute, only a few times a week [27].
Figure 1: Mars missions timeline with the data rates of their links to Earth.
To build high-capacity DS communications, research efforts have been recently devoted to exploring the potentials of free space optical (FSO) communications [28, 13, 29, 30, 31, 32]. Compared to radio frequency (RF), FSO links enjoy larger bandwidth, smaller apertures at final stations, much improved security, less transmit power, and higher immunity against interference. Although most recent DS Mars missions are RF-based [16], many projects have already adopted FSO for Moon and Mars communications with plans to expand to other planets such as Jupiter [33, 34, 35, 36, 37]. In addition, several implementation-based studies have been done to evaluate the technology transition, e.g. [30, 38]. Tremendous work is yet to be done though to make FSO technology ready for future missions [39].
Adopting FSO in DS will play an important role in implementing interplanetary broadband networks. However, achieving such a goal will depend on the reliability, effectiveness, and efficiency of many other technologies, including modulation, coding, detection, and networking protocols. The criteria for selecting such schemes differ tremendously from those for terrestrial, airborne, and even near-Earth communications. In this survey, we highlight the various DS optical communications (DSOC) design criteria and present the latest research and innovations in the area.
@@ I-B Related Surveys
There are few surveys available on DS communications [29, 31, 36, 40, 30, 41, 42, 43, 44], most of which are outdated and/or do not address the field comprehensively. Although [29] covers various aspects of DSOC, an updated survey that reflects the current field demands, challenges, and advances is needed. The survey of [31] covers some aspects of optical communications in DS, but only partially. Biswas et al. [36] focus merely on the progress made and the status of DSOC in National Aeronautics and Space Administration (NASA) projects. In the same direction, [40] overviews mainly laser-based DS demonstration projects and future missions. Meanwhile, [30] provides only studies about DS systems architects and discussions about the transition from RF to optical technology. In addition, [41] focuses only on DS networking aspects. Similarly, [42] covers merely several aspects of RF-based DS communications systems. Furthermore, [43] addresses FSO and RF challenges and discusses DS systems implementations and demonstrations; however, it does not present research discussions and future perspectives in the field. Finally, [44] gives an inspiring and relatively recent overview of DS networking and implementations, yet, it lacks any further details in the other aspects, including physical characteristics and field advances.
Despite their valuable contributions, none of the above works covers DS communications holistically. In addition, most of these surveys are not recent. In contrast, as we demonstrate in Table I, our work aims to cover the key aspects of DSOC, add the missing areas from the aforementioned related work, and discuss the latest advances in the field, including orbital angular momentum (OAM), interplanetary network (IPN), quantum communications (QC), and the RF/FSO which is an integration of both RF and FSO technologies.
@@ I-C Contributions
In this survey, we cover existing DSOC systems’ limitations, recommended technologies to cope with them, and challenges facing their adoption. The contributions of this paper relative to the literature are summarized as follows.
@@
    We cover the major challenges of DS communications. We also discuss DSOC promises, challenges, and potential solutions and compare them with the existing RF technology.
@@
    We survey the widely used modulation and coding schemes and the opportunities for enhancing performance by fusing them with emerging detection techniques.
@@
    We review recommended networking architecture models, present key characteristics of promising protocols, and identify related open research directions.
@@
    We present an overview of the IPN and study the visions for its future.
@@
    To the best of our knowledge, we provide the first thoughtful discussion about implementing OAM for DS and identify challenges and potential solutions.
@@
    We discuss the advantages, efforts, and challenges of implementing QC and partnering RF and FSO technologies in DS.
@@
TABLE I: COMPARISON OF DS COMMUNICATIONS SURVEYS
@@
Overview Missions timeline $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Demonstrations $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Existing systems $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Limitations $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Implementations $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Future perspectives $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Basics Link engineering $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
Overall challenges $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
Enabling technologies Networking protocols $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Modulation $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
Coding $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
Receiver technologies $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Performance analysis Capacity $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
Advances OAM $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
IPN $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
QC $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
RF/FSO $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$
@@ I-D Organization
The remainder of the paper is structured as follows. Section II outlines DSOC advantages over RF and implementation efforts and challenges. Section III analyzes DS links physical characteristics and discusses related research directions. While Section IV characterizes DSOC capacity limits, Section V surveys DS networking protocols and related research efforts. Section VI investigates on the state-of-the-art and challenges of DSOC advances including, IPN, OAM, QC, and RF/FSO. Finally, Section VII concludes the survey.
@@ II DSOC Advantages and Challenges
Optical communications promise tens to hundreds of times increase in data rates, thanks to the high-available bandwidth in the license-free spectrum in which the technology operates [31, 45]. Such a capacity improvement has already been proven in demonstrations [46]. To illustrate further the gain of using optical communications compared to RF in DS, based on [39], the figure of merit (FoM) metric can be defined as follows:
@@
where $D$ is the end-to-end DS link distance in the astronomical unit (AU). Here, $R$, $P_{t}$, and $D_{r}$ are the data rate in bits per second, transmit power in Watts, and receiver diameter size in meters, respectively. As it accounts for a space terminal performance measured through data rate and cost reflected through receiver aperture size and communications power, the FoM metric provides a direct comparison between the RF and FSO technologies at the system level at any distance.
Based on the projections available in [14, 40, 46], Fig. 2 demonstrates FoM for Lunar and Mars. Unsurprisingly, a Lunar FSO system represents more than 15 dB gain over its RF counterpart. The improvement seems to be even more substantial over Mars distances. However, implementing FSO in DS communications faces a number of unique challenges.
Below, we discuss different challenges facing DS communications at the link level and the DSOC link budget along with the challenges in optical communications.
@@ II-A DS Communications Challenges
Ensuring high-capacity and reliable communications in DS faces the challenges listed below.
  • Ultra-long, which based on the International Telecommunications Union exceeds 2 million kilometers, and time-changing distances between Earth and probes (satellites in other planets’ orbiters, rovers, etc.).
@@
The communications with targets in DS require sending signals over very long distances, leading to strong power dissipation and hence huge signal attenuation. Even with relatively large-gain transmit and receive antennas, the received power can still be very weak. Also, the long-distance links result in an inevitable significant latency that prevents real time communications and control. For example, a two-way communications with Mars may undergo 40 minutes transmission delay.
Furthermore, the Solar flares are shown to make communications signals weaker and noisier [29]. Besides, Solar scintillation and turbulence lead to performance degradation. It is important to carefully account for the near-Sun effect on the signal strength to compensate for any losses and prevent any damage to probe systems’ circuits. Also, due to the movement of Earth and probes, the Sun may completely block communications path.
In addition, SWaP constraints are of major concerns not only for efficient spacecraft design but also for communications performance improvement. Designing a space mission with ultra-lightweight resources is thought of as a way to utilize a better communications performance [29]. There are technologies such as Gossamer that has the premise to design spacecraft structures with thin, low-modulus, and low-mass materials. An example of Gossamer technology application is the inflatable antennas; which could have an areal density of a fraction of a kilogram per square meter [47]. Also, in the direction of optical communications, the development of ultra-lightweight telescopes is realized; an example of which is the Cassegrain telescope which has composite mirrors with areal density in the range of 1–10 kg/m2 [48].
Lastly, existing terrestrial technologies, especially at the physical and protocol levels, can not be adopted directly in DS. For example, as they are simpler in implementation, less lossy, and more power and bandwidth-efficient, digital modulation schemes are preferred in DS RF-based communications. Nevertheless, amplitude-based modulations that can not be generated with nonlinear amplifiers are not suitable for DS as working in saturation mode is needed to maximize power efficiency [49]. Also, in networking, commercial terrestrial network architectures, for example transmission control protocol (TCP)/internet protocol (IP), are not efficient to ensure data delivery in the DS time changing environments [50].